WAC Notes February 17 2020

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes CREX << Instructions

WAC Notes February 16 2020 << >> WAC Notes February 18 2020

February 17 2020

Daily Meeting BlueJeans


Production Run List - Slug 136 - IHWP IN - Flip-RIGHT - Production

Start this slug at ~10:04am, February 16

  • Good - Run 6439
    • Aq slightly larger in minirun 0, detector width is also large. USR has ~11% residual correlation with AqDS
    • ~10 ppm larger detector slope
  • Good - Run 6440
    • LQ1 tripped, applied event_cut at the end of the run
    • Detector width is still large
    • BCM asym noisy in the first half
  • Good - Run 6441
    • LQ1 tripped, applied event_cut at the end of the run
    • Very short run
    • Detector width is still large
    • Large residual correlation of detector (USR) with AqUS and AqDS
  • Good - Run 6443
    • Detector width is still large
  • Good - Run 6444
    • Detector width is still large
  • Good - Run 6445
    • Large Aq for minirun 4 (4.88+/-2.30)
    • Detector width is still large
  • Good - Run 6446
    • Energy BPM wires saturated
    • Detector width is still large
  • Good - Run 6447
    • Energy BPMs saturated many times. FFB ON/OFF to investigate this issue
  • Good - Run 6448
    • Some beam excursion in the beginning of the run
  • Good - Run 6449
    • ~ 1mm of beam excursion at the end. Quite large Aq in 0th and 2nd minirun
  • Good - Run 6451
    • ~ 1mm beam excursion in the beginning
  • Good - Run 6452
    • ~0.8mm of beam excursion at the end. 4ec again sees this huge. bpm12X also has ~0.8mm-1mm of excursion
    • Due to the excursion detector yield has changed (increased in USR and decrease in USL). SAMMs also see this.
  • Good - Run 6455
  • Good - Run 6456
  • Good - Run 6457
  • Good - Run 6458
    • Aq little bit noisy at the end part of the run
  • Good - Run 6459
    • Some intermittent energy excursions ~1mm

Production Run List - Slug 137 - IHWP OUT - Flip-RIGHT - Production

Start this slug at ~04:02am, February 17

  • Good - Run 6460
    • An outlier in MAIN det asym
  • Good - Run 6461
    • Some beam excursions but looks good
  • Good - Run 6462
  • Good - Run 6463
    • Large Aq in 0th minirun (~4.58+/-1.88 ppm)
    • Don't see any obvious reason for this
  • Good - Run 6464

Non-Production Runs and Bad Runs

  • run6442 --> Junk
  • run6450 --> Junk (Carbon 1%, MCC trying to fix energy drift)
  • run6453 --> Junk (Lots of GAIN search failure)
  • run6454 --> Junk (Carbon 1%, MCC trying to fix FFB)

Pedestal/Calibrations Updates

Issues

Updates